OpenStudy (anonymous):

One half of Ron's age plus one third of Bob's age equals 6. Bob's age is two more than one half of Ron's age. How old is Ron? How old is Bob? (use equations!)

zepdrix (zepdrix):

R <-- Ron's age B <-- Bob's age (1/2)B + (1/3)B = 6 <-- half Ron's age + one third Bob's age equals 6. (1/2)R <-- one half Ron's age (1/2)R + 2 <-- 2 more than one half Ron's age (1/2)R + 2 = B <-- Bob's age is 2 more than one half Ron's age So we have a system of equations. (1/2)B + (1/3)B = 6 (1/2)R + 2 = B You can use substitution or elimination to solve. If you're still confused, let me know! :)

OpenStudy (anonymous):

ok thank you, please five me a minute to try to solve!

OpenStudy (anonymous):

give*

OpenStudy (anonymous):

isnt it (1/2)B in the first equation?

OpenStudy (anonymous):

i mean (1/2)R

zepdrix (zepdrix):

i wrote B twice didn't i? lol silly me XD

zepdrix (zepdrix):

(1/2)R + (1/3)B = 6 Yah sorry about that c:

OpenStudy (anonymous):

lol its ok ok please give me like 5-10 minues to solve and I'll tell you what i got!

OpenStudy (anonymous):

i got really weird answers :( i think i did it wrong. can you help me please?

OpenStudy (anonymous):

i got r= (8/9) and b=2.444

OpenStudy (anonymous):

thats not right?

zepdrix (zepdrix):

Hmm I keep getting different answers depending on the way I solve this :D lemme go slowly through it and see where we're getting stuck.

OpenStudy (anonymous):

ok thanks please explain it to me because im stuck

zepdrix (zepdrix):

 I really dislike working with fractions, it's good to get them out of the way immediately if you can :) Let's try elimination from here.